OpenStudy (jchick):

Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work.

OpenStudy (jchick):

I know the slope intercept form answer but not how to explain it.

OpenStudy (jchick):

In part A I had to find the Point Slope Form this is y + 2 = 2(x - 3)

OpenStudy (jchick):

The slope intercept form is y = 2x - 8

OpenStudy (jchick):

@pooja195 please help I don't know how to explain how I got the answer

pooja195 (pooja195):

ok well you have to solve this \[ y + 2 = 2(x - 3) \]

pooja195 (pooja195):

when you do you get y = 2x - 8 get it?

OpenStudy (jchick):

Yeah I do I just need to show work

OpenStudy (jchick):

So could you tell me if my work looks right

OpenStudy (jchick):

subtract two from each side Solve for y: y+2 = 2 (x-3) Solve for y. Subtract 2 from both sides: Answer: y = 2x - 8

OpenStudy (jchick):

@pooja195